What is the pH of 0.15M of hydroxylamine of kb=1.1*10^-8
1
Expert's answer
2020-10-05T14:21:45-0400

Solution:

Hydroxylamine, NH2OH, is a weak base.

The equilibrium equation for its reaction with water:

NH2OH(aq) + H2O(l) ⇌ NH3OH+(aq) + OH(aq);  Kb = 1.1×10-8


From this, we can construct an ICE table:



The Kb expression for NH2OH is:



Kb = [ x ] × [ x ] / [ 0.15 - x ]


C(NH2OH) / Kb = (0.15) / (1.1×10-8) = 1.364×107 > 500

Since the ratio is greater than 500, we can remove (–x) from the equation.


The equation simplifies to:

Kb = [ x ] × [ x ] / [ 0.15 ]

Kb = x2 / 0.15

Solving for x:

x2 = 1.1×10-8 × 0.15 = 0.165×10-8

x = (0.165×10-8)0.5 = 4.062×10-5


This means [OH] = [NH3OH+] = x = 4.062×10-5 M. 


From [OH], we can calculate pOH:

pOH = -log[OH-] = -log(4.062×10-5) = 4.39


From pOH, we can calculate pH:

pH + pOH = 14;

pH = 14 - pOH = 14 - 4.39 = 9.61

pH = 9.61


Answer: pH = 9.61